KpopAsianDramaFanYT:

The volume of a rectangular prism is represented by the function x^3 + 11x^2 + 20x − 32. The width of the box is x − 1, while the height is x + 8. Find the expression representing the length of the box.

jhonyy9:

do you know formula of volume for a rectangular prism ?

jhonyy9:

V = area of base time height

jhonyy9:

formula for area of a rectangle ?

KpopAsianDramaFanYT:

ahh ok

jhonyy9:

can you solve it now ?

jhonyy9:

@Laylalyssa

KpopAsianDramaFanYT:

no

jhonyy9:

so than what mean this above wrote ,,ok" ?

KpopAsianDramaFanYT:

nvm that

KpopAsianDramaFanYT:

please help me

jhonyy9:

so to get the expression of length you need divide the expression of volume by the product of expressions of height and width ok ?

jhonyy9:

multiplie the x-1 by x +8 and what you get so divide the expression of volume by this expression what you get

jhonyy9:

@supie

KpopAsianDramaFanYT:

ok

jhonyy9:

questions ?

KpopAsianDramaFanYT:

nope

jhonyy9:

ok hnd cya

KpopAsianDramaFanYT:

x^3 + 11x^2 + 20x − 32 divided by 8?

jhonyy9:

(x^3 + 11x^2 + 20x − 32) divided by [(x − 1)( x + 8)] = ?

jhonyy9:

first multiplie (x-1)(x+8)= ?

jhonyy9:

= x^2 +8x -x -8 = x^2 +7x -8 and now just you need divide the volume expression by this x^2 +7x -8 and so will get the expression of the length

jhonyy9:

do you can ?

jhonyy9:

(x^3 + 11x^2 + 20x − 32) : (x^2 +7x -8) = ?

jhonyy9:

do you know the way ?

KpopAsianDramaFanYT:

is it x - 4?